A species of fish grows 1.2 cm in first year. Each year it grows by 2% more than it did the year before. Using geometric sequence find, how much does it grow in 10th year.



1
Expert's answer
2022-06-13T16:13:44-0400

Each year the growth is 2%=0.02 more than the previous year, so we have the geometric secuence with the first term 1.2 and the common ratio "r=1+0.02=1.02."

We know that for geometric secuence "a_n=a_1\times r^{n-1},"

so in 10th year this species of fish grows

"a_{10}=1.2\times 1.02^9=1.4341110823467733 \approx 1.43" cm.


Answer: 1.43 cm